OpenStudy (anonymous):

A cup of yogurt cost $2.60 plus ten cents per topping. A cup of ice cream costs $2.00 plus twenty five cents per topping. How many toppings are needed so that the two desserts cost the same?

OpenStudy (amistre64):

when does .10x+2.60 = .25x + 2.00 ?

OpenStudy (amistre64):

.10x+2.60 = .25x + 2.00 -.10x -.10x ------------------------ 2.60 = .15x +2.00 -2.00 -2.00 --------------------- .60 = .15x ; solve for x